OpenStudy (anonymous):

Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y). Part B: How much time does Ted spend on practicing the breast stroke every day? Part C: Is it possible for Ted to have spent 45 minutes practicing the butterfly stroke every day?

OpenStudy (texaschic101):

butterfly stroke = x breast stroke = y x + y = 50 y = x + 20 these are your two equations now sub x + 20 in for y in the other equation x + y = 50 x + x + 20 = 50 2x + 20 = 50 2x = 50 - 20 2x = 30 x = 15 now sub 15 in for x in either of the equations to find y x + y = 50 15 + y = 50 y = 50 - 15 y = 35 He spends 15 minutes on butterfly stroke (x) He spends 35 minutes on breast stroke (y) It is not possible for him to spend 45 minutes on butterfly stroke, unless he only spends 5 minutes on breast stroke. He is only there 50 minutes.

OpenStudy (texaschic101):

any questions ?